The 4.5 percent bond of JL Motors has a face value of $1,000, a maturity of 7 years, semiannual interest payments, and a yield to maturity of 6.23 percent. What is the current market price of the bond?

$945.08

$947.21

$903.05

$959.60

$912.40
Answer #1
1 Face value (FV) $                                         1,000
2 Coupon rate 4.50%
3 Number of compounding periods per year 2
4 = 1*2/3 Interest per period (PMT) $                                         22.50
5 Number of years to maturity 7
6 = 3*5 Number of compounding periods till maturity (NPER) 14
7 Market rate of return/Required rate of return 6.23%
8 = 7/3 Market rate of return/Required rate of return per period (RATE) 3.12%
Bond price PV(RATE,NPER,PMT,FV)*-1
Bond price $                                       903.05